OpenStudy (anonymous):

A graphic designer and her assistant working together can complete an advertising layout in 6 days. The assistant working alone could complete the job in 16 more days than the designer working alone. How long would it take each person to do the job?

OpenStudy (amistre64):

so the amount of time it takes the designer to do a job is 1/1 the assisstant will do 1/17 of it in the same amount of time, right?

OpenStudy (amistre64):

if the job done together is done in 6 days that means the designer did 6 parts to the assistants 6/17 parts

OpenStudy (anonymous):

where did 1/17 come from?

OpenStudy (anonymous):

i'm supposed to declare variables and have equations and pellet

OpenStudy (anonymous):

and pellet*

OpenStudy (amistre64):

its the information in the posted problem; it takes 16 more days to finish the same amount of work; D = 1 day done A = 1 + 16 days done

OpenStudy (amistre64):

the ratio is 1 to 17 then; right?

OpenStudy (anonymous):

its takes the assistant alone 16 more days than it would take the designer working alone..

OpenStudy (amistre64):

i see, i read it wring :)

OpenStudy (anonymous):

lol okay! you were confusing me for a minute!

OpenStudy (amistre64):

The assistant working alone could complete the job in 16 more days than the designer working alone. D = D days A = D+16 days add them together and we get: 2D + 16 days to do 2 of the same jobs right? im thinking here

OpenStudy (amistre64):

2D + 16 would be done in 12 days then ... that make sense?

OpenStudy (anonymous):

D = # of days designer could do it alone A = # of days Assistant could do it alone ?

OpenStudy (amistre64):

yes, it was either that or come up with something that would just confuse the situation :)

OpenStudy (anonymous):

okay; so one equation would be: a=2d+16 what would the other one be?

OpenStudy (anonymous):

a+d=6 ?

OpenStudy (amistre64):

A+D = 6, yes if we knew what A and D where .... but all we know is what it is if they workd alone. right?

OpenStudy (anonymous):

but when they work together; they can complete it in 6 days.

OpenStudy (amistre64):

right, so lets assume they do 2 of the same jobs; they would finish in 12 days correct?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

So if D can finish in D days on one job; let her do that job alone A can finish the other job in D+16 days, so let A do the other one .... is my thinking

OpenStudy (amistre64):

D + (D+16) = 12 days for 2 jobs then, and this might be more work then needed, but to me it makes sense

OpenStudy (anonymous):

i'm not following

OpenStudy (amistre64):

let me try to see if this is going to work first, then ill try to describe it if it does :)

OpenStudy (anonymous):

haha okay.

OpenStudy (amistre64):

2d + 16 = 12 days 2d+16 ------ = 1 days work ; reduce 12 d + 8 ----- = 1 days work 6 d/6 + 8/6 maybe, and i cant be sure yet, the A does the same amount of wok in 1' 1/3 the time it takes D to do it

OpenStudy (amistre64):

6/8 = .75, of the work the D does .... .25 + .75 = 1, and thats done in 6 days; but how does this measure against the rest of it?

OpenStudy (amistre64):

i wish this made more sense to me than a gut feeling :)

OpenStudy (anonymous):

2d+16=12 : where is that coming from?

OpenStudy (amistre64):

D=8 days; then A = 24 days if im gutting it right 1/8 + 1/24 = 4/24; 24/4 = 6 all right

OpenStudy (amistre64):



OpenStudy (amistre64):

spose we have 2 jobs to do in 12 days; since 6 days * 2 = 12

OpenStudy (anonymous):

why not just a=d+16 though?

OpenStudy (amistre64):

because to me, that tells us nothing in the outcome of things; that is just a relative measure and nothing that can be quantified with a single job. We have to set up an experiment to measure against.

OpenStudy (amistre64):

in other words, that is related to just this one job and we cant determine, or at least i cant determine what to measure it against ..

OpenStudy (amistre64):

we know that alone it will take 16 more days to finish alone, than it does with help. we know that D can finish it alone in ..... some time frame that is undetermined

OpenStudy (amistre64):

Without being able to establish the timeframe for D how do we measure another 16 days from it?

OpenStudy (amistre64):

give D and A their own workspace to measure against; 2 jobs that would have taken them 12 days to complete if worked together, and D works one, while A works the other for comparison

OpenStudy (amistre64):

Job = D = 6days Job = D+16 = 6days 2Jobs = 2D+16 = 12days; reduce by 2 Job = D + 8 = 6 days; divide by 6 to see how much is done in 1 day 1/6 Job = D/6 + 8/6 = 1 day .... we are getting somewhere now

OpenStudy (amistre64):

these are percentages of work done in the day; 6/D and 6/8 are the speed with which the work relative to each other, and the faster one HAS to be the Designer. 6/8 = 3/4 = .75 = 75%; leaving 25% for the Assistant

OpenStudy (amistre64):

the rest of this is hard to put into words ... but im trying :)

OpenStudy (anonymous):

I think you are doing a great job!!

OpenStudy (amistre64):

... thnx :)

OpenStudy (amistre64):

\[\frac{6}{D}:\frac{6}{8},\ D=8\] but why?

OpenStudy (amistre64):

would it be the same if the length was 5 extra days? 2d+5 = 12 (2d+5)/12 = 1 12/2d = 12/5; d = 5/2 = 2.5 A = 7.5 6(1/2.5 + 1/7.5) = 3.2 ..... so it seems to be a fluke

OpenStudy (amistre64):

thats the best i got fer tonite, good luck with it :)

OpenStudy (anonymous):

How does this work? A=16+G; G=graphic designer and A=assistant A+G=6; so we can plug A in there (16+G)+G=6 16+2G=6 2G=-10 G=-5, hmmm I dont think we can have a neg amount of days, we could plug this in to find A A=16+(-5) A=11

OpenStudy (anonymous):

I don't know, it was just a thought :)

OpenStudy (anonymous):

Here: Designer working alone takes t days Assistant takes t +16 days Together they work for 6 days; \[6/t+6/(t+16)=1\] clear the denominators and combine terms: \[6(t+16) + 6t=t(t+16)\] \[t^2+4t-96=0\] \[(t+12)(t-8)=0\] t must be 8 (positive time to complete job for designer Designer takes 8 days Assistant takes 24 days Together they take 6 \[6/8+6/24=1\]

OpenStudy (anonymous):

Solve the following for x, the time required for the graphic designer to complete the layout solo.\[\frac{1}{x+16}+\frac{1}{x}= \frac{1}{6}\]x=8, x+16 = 24